Sir Marvin believes that the mean score of students who attend his review class in

Mathematics is 35. Forty five students were randomly selected and gathered their scores. It

is revealed that the sample mean is 38 and the standard deviation is 8.4 points. Using 0.01

level of significance, test the hypothesis.


1
Expert's answer
2021-05-07T09:44:36-0400

M = 38

Z = 2.58

sM = "\sqrt{(\frac{8.4^2}{45})}" = 1.25


μ = M ± Z(sM)

μ = 38 ± 2.58"\times"1.25

μ = 38 ± 3.23


Answer: sir Marvin's hypothesis is correct at 0.01 level of significance.